# Flow Snapshot Neurons in Action: Deep Neural Networks Generalize to Biological Motion Perception

### v1

@@ Abstract
Biological motion perception (BMP) refers to humans’ ability to perceive and recognize the actions of living beings solely from their motion patterns, sometimes as minimal as those depicted on point-light displays. While humans excel at these tasks without any prior training, current AI models struggle with poor generalization performance. To close this research gap, we propose the Motion Perceiver (MP). MP solely relies on patch-level optical flows from video clips as inputs. During training, it learns prototypical flow snapshots through a competitive binding mechanism and integrates invariant motion representations to predict action labels for the given video. During inference, we evaluate the generalization ability of all AI models and humans on 62,656 video stimuli spanning 24 BMP conditions using point-light displays in neuroscience. Remarkably, MP outperforms all existing AI models with a maximum improvement of 29% in top-1 action recognition accuracy on these conditions. Moreover, we benchmark all AI models in point-light displays of two standard video datasets in computer vision. MP also demonstrates superior performance in these cases. More interestingly, via psychophysics experiments, we found that MP recognizes biological movements in a way that aligns with human behavioural data. All data and code will be made public.
@@ 1 Introduction
Figure 1: Humans excel at biological motion perception (BMP) tasks with zero training, while current AI models struggle with poor generalization performance. AI models are trained to learn to recognize actions from natural RGB videos and tested using BMP stimuli on point-light displays, which come in two forms: Joint videos, which display only the detected joints of actors in white dots, and Sequential position action videos (SP), where light points in white are randomly positioned between joints and reallocated to other random positions on the limb in subsequent frames (Sec. 3.1). Note that skeletons, shown in gray in the example video, are not visible to humans or AI models during testing. The generalization performance of both humans and models is assessed after varying five properties in temporal and visual dimensions. Example videos are available at: link.
@@
In parallel to the studies of BMP in psychology and neuroscience, action recognition on images and videos in computer vision has evolved significantly over the past few decades due to its wide range of real-world applications [15, 37, 75, 76, 50, 67]. The field has progressed from relying heavily on hand-crafted features [91, 68, 47] to employing sophisticated deep-learning architectures  [87, 11, 93, 27, 29, 28, 24, 4] capable of capturing the temporal dynamics and spatial configurations of complex activities within dynamic, unstructured environments  [74, 85]. Despite these advancements, existing AI models still struggle with generalization issues related to occlusion [95, 3, 51], noisy environments  [100], viewpoint variability  [26, 54, 92], and subtle human movements  [77, 39]. While various solutions have been proposed to enhance AI generalization in action recognition [12, 65, 14, 63, 96, 42, 52, 102, 73, 43], they do not specifically tackle the generalization challenges in BMP tasks.
@@
Subsequently, to enhance the generalization capability of AI models, we draw inspiration from  [30] in neuroscience and introduce Motion Perceiver (MP). MP only takes dense optical flows between any pairs of video frames as inputs and predicts action labels for the given video. In contrast to many existing pixel-level optical flow models [83, 84, 79, 91, 71], MP calculates dense optical flows at the granularity of patches from the feature maps. In MP, we introduce a set of flow snapshot neurons that learn to recognize and store prototypical motion patterns by competing with one another and binding with dense flows. This process ensures that similar movements activate corresponding snapshot neurons, promoting consistency in motion pattern recognition across patches of video frames. The temporal dynamics within dense flows can vary significantly depending on factors such as the speed, timing, and duration of the actions depicted in video clips. Thus, we also introduce motion-invariant neurons. These neurons decompose motions along four orthogonal motion directions and integrate their magnitudes over time. This process ensures that features extracted from dense flows remain invariant to small changes and distortions in temporal sequences.
@@ 2.1 Patch-level Optical Flow
Pixel-level optical flow has been a common approach to modelling temporal dynamics in videos  [83, 84, 79, 91, 71]. Different from these works, we use the frozen ViT  [20], pre-trained on ImageNet  [18] with DINO [10], to extract feature map $F_{t}\in\mathbb{R}^{N\times C}$ from $I_{t}$ and compute its optical flow relative to other frames. Considering $F_{t}$ as a 2D grid of patches in $N=H\times W$ where $H$ and $W$ are the height and width of $F_{t}$. $N$ represents the number of patches in the feature map and $C$ is the feature dimension per patch. A unique spatial location for each patch can be defined by its 2D coordinates. We concatenate the X and Y coordinates of all patches in $F_{t}$ to form the patch locations $G\in\mathbb{R}^{N\times 2}$.
@@
Next, we introduce how the patch-level optical flow for video $\mathcal{V}$ is computed. Without loss of generality, given any pair of feature sets $a\in\mathbb{R}^{m\times d}$ and $b\in\mathbb{R}^{n\times d}$, where $m$ and $n$ are the numbers of patches in the feature sets and $d$ is the feature dimension, the adjacency matrix $Q_{(a,b)}$ between $a$ and $b$ can be calculated as their normalized pairwise feature similarities: $Q_{(a,b)}=\frac{e^{f(a)f(b)^{T}/\tau}}{\sum\limits_{n}e^{f(a)f(b)^{T}/\tau}}\in\mathbb{R}^{m\times n}$, where the superscript $T$ is the transpose function, $f(\cdot)$ is the $l_{2}$-normalization, and $\tau$ is the temperature controlling the sharpness of distribution with its smaller values indicating sharper distribution. We set temperature $\tau=0.001$ and the influence of $\tau$ is analyzed in Tab. 4.
@@ 2.2 Flow Snapshot Neurons
Drawing on findings from neuroscience that highlight biological neurons selective for complex optic flow patterns associated with specific movement sequences in motion pathways  [30], we introduce “flow snapshot neurons." These neurons are designed to capture and represent prototypical moments or “slots" in movement sequences. Mathematically, we define $K=6$ flow snapshot neurons or slots $\hat{Z}\in\mathbb{R}^{K\times D}$, where $D=2\times(T-1)$ is the feature dimension per slot. $\hat{Z}$ contains learnable parameters randomly initialized with the Xavier uniform distribution  [32]. The impact of $K$ is discussed in Tab. 4.
Slot attention mechanism  [56] separates and organizes different elements of an input into a fixed number of learned prototypical representations or “slots." The learned slots have been useful for semantic segmentation [94, 53, 101, 44, 21]. Here, we follow the training paradigm and implementations of the slot attention mechanism in  [56] and apply it to $\hat{O}$. Formally, based on the cross-attention mechanism [90], slots $\hat{Z}\in\mathbb{R}^{K\times D}$ serve as the query, $\hat{O}\in\mathbb{R}^{S\times D}$ where $S=T\times N$ and $D=2\times(T-1)$ contribute to both the key and the value, and $q(\cdot),k(\cdot),v(\cdot)$ are the linear transformation employed to project inputs and slots into a common dimension $B$, which can be formulated as:
$\text{attn}_{i,j}:=\frac{e^{J_{i,j}}}{\sum_{l}e^{J_{i,l}}}\quad\text{ where }\quad J:=\frac{1}{\sqrt{D}}k(\hat{O})\cdot q(\hat{Z})^{T}\in\mathbb{R}^{S\times K},$ (2)
$\hat{Z}_{p+1}=\text{GRU}\left(\hat{Z}_{p},h_{p}\right).$ (4)
$L_{slot}=\text{CE}(Q_{(\hat{Z},\hat{O})}Q_{(\hat{O},\hat{Z})},\mathbf{I}),$ (5)
where $Q_{(a,b)}$ is defined as the normalized pairwise feature similarities between $a$ and $b$ in Sec. 2.1. We use the temperature $\mu$ of 0.05 in $Q_{(\hat{Z},\hat{O})}$ and $Q_{(\hat{O},\hat{Z})}$. CE($\cdot$,$\cdot$) stands for the cross-entropy loss and $\mathbf{I}\in\mathbb{R}^{K\times K}$ is the identity matrix. The effect of $\mu$ is covered in Tab. 4. During inference, we keep $\hat{Z}$ frozen and leverage the activations of flow snapshot neurons for action recognition, denoted as $\hat{M}=f(\hat{O})f(\hat{Z})^{T}$, where $\hat{M}\in\mathbb{R}^{T\times N\times K}$.
Unlike  [99] that use slot-binded optical flow sequences $\hat{M}\hat{Z}\in\mathbb{R}^{T\times N\times D}$ for downstream tasks, we highlight two advantages of using $\hat{M}$. First, the dimensionality of flow similarities ($K$) is typically much smaller than that of slot-binded optical flow sequences ($D$), effectively addressing overfitting concerns and reducing computational overhead. Second, by leveraging flow similarities, we benefit from slot activations that encode prototypical flow patterns distilled from raw optical flow sequences. This filtration process helps eliminate noise and irrelevant information from the slot-binded flow sequences, enhancing the model’s robustness.
@@ 2.3 Motion Invariant Neurons
The temporal dynamics in video clips can vary significantly in the speed and temporal order of the actions. Here, we introduce motion-invariant neurons that capture the accumulative motion magnitudes independent of frame orders. Specifically, we define the optical flow $(O^{x,n,I_{t}}_{m\rightarrow t},O^{y,n,I_{t}}_{m\rightarrow t})$ for patch $n$ in $I_{m}$ moving from frame $I_{m}$ to $I_{t}$ along $x$ and $y$ axes. Every patch-level optical flow in $\hat{O}$ can be projected into four orthogonal motion components along $+x$, $-x$, $+y$ and $-y$ axes. Without loss of generality, we show that for patch $n$ in $I_{t}$, all its optical flow motions over $T$ frames along the $+x$ axis are aggregated:
@@
By repeating the same operations along $-x,+y,-y$ axes for all the $N$ patches over $T$ frames, we obtain motion invariant matrix $\tilde{M}\in\mathbb{R}^{T\times N\times 4}$. A stack of self-attention blocks followed by a global average pooling layer fuse information along 4 motion components and then along temporal dimension $T$. To encourage the model to learn motion invariant features for action recognition, we introduce cross-entropy loss $L_{invar}$ to supervise the predicted action against the ground truth $\mathcal{Y}$.
@@ 2.4 Multi-scale Feature Fusion and Training
While $\hat{M}_{1,2,4,8}$ captures the detailed temporal dynamics, $\tilde{M}$ learns robust features against variations in temporal orders. MP combines feature vectors from $\hat{M}_{1,2,4,8}$ and $\tilde{M}$ with a fully connected layer and outputs the final action label. A cross-entropy loss $L_{fuse}$ is used to balance the contributions from $\hat{M}_{1,2,4,8}$ and $\tilde{M}$. The overall loss is: $L=\alpha L_{slot}+L_{flow}+L_{invar}+L_{fuse}$, where the loss weight $\alpha=10$. See Tab. 4 for the effect of $\alpha$.
Implementation Details. Our model is trained on Nvidia RTX A5000 and A6000 GPUs, and optimized by AdamW optimizer [58] with cosine annealing scheduler [57] starting from the initial learning rate $10^{-4}$. Data loading is expedited by FFCV [49]. All videos are downsampled to $T=32$ frames. Our model is trained on the training sets of RGB videos in BMP, NTU RGBD+60 and NW-UCLA datasets respectively. BMP: Videos are resized to $224\times 398$ pixels for 50 epochs training with a batch size of 48. NTU RGB+D 60: Videos are resized to $224\times 398$ pixels and trained for 100 epochs with a batch size of 128. NW-UCLA: Videos are resized to $224\times 300$ pixels and trained for 100 epochs using a batch size of 16. We use a random crop of $224\times 224$ pixels with horizontal flip for training, and a central crop of $224\times 224$ pixels for inference. We use the same set of hyper-parameters for our model in all the datasets.
@@ 3.1 Our Biological Motion Perception (BMP) Dataset with Human Behavioral Data
We introduce the BMP dataset, comprising 10 action classes: pick up, throw, sit down, stand up, kick something, jump up, point to something, nod head/bow, falling down and arm circles. They are specifically chosen for their strong temporal dynamics and minimal reliance on visual cues. This selection criterion ensures that specific objects or scene contexts, such as a soccer ball on green grass, do not bias the AI models toward recognizing the action as “playing soccer.” This approach aligns with our research focus on generalization in motion perception. In the BMP dataset, there are three types of visual stimuli.
@@
Temporal order (TO): To disrupt the temporal order  [78], we reverse (Reversal) or randomly shuffle (Shuffle) video frames for two types of videos, RGB and Joint above.
Temporal resolution (TR): We alter the temporal resolution  [78] by uniformly downsampling 32 frames to 4 or 3 frames, labelled as 4-frames and 3-frames, respectively. For models that need a fixed frame count, each downsampled frame is replicated multiple times before advancing to the next frame in the sequence, until we reach the necessary quantity.
Amount of visual information (AVI): We quantify the amount of visual information based on the number of light points (P) in Joint videos  [40]. Specifically, we included conditions: 5, 6, 10, 14, 18, and 26 light points. To generate Joint videos, Alphapose [25] is used as the tool to detect the joints of a human body from RGB videos. After that, we use these detected joints to generate BMP stimulus on point-light displays. The specific positions of light points correspond to certain joint locations:
  • J-5P: 1 point each on the nose, 1 point each on one hand and one ankle.
Lifetime of visual information (LVI): In SP videos, we manipulate the number of consecutive frames during which each point remains at a specific limb position before being randomly reassigned to a different position. Following [6], we refer to this parameter as the “lifetime of visual information" (LT), and include LT values of 1, 2, and 4. A longer lifetime implies that each dot remains at the same limb position for a longer duration, thereby conveying more local image motion information.
Invariance to camera views (ICV): Neuroscience research  [38] has revealed brain activity linked to decoding both within-view and cross-view action recognition. To probe view-dependent generalization effects, we sorted video clips in J-6P into three categories based on camera views: frontal, $45^{\circ}$, and $90^{\circ}$ views, which are respectively labelled as J-6P-0V, J-6P-45V and J-6P-90V in short form. $45\degree$ view and $90\degree$ view are rotated either clockwise or counterclockwise from the frontal view.
Human psychophysics experiments: We conducted human psychophysics experiments schematically illustrated in Fig. 3, using Amazon Mechanical Turk (MTurk)  [8]. A total of 60 subjects were recruited. The experiment consists of 13,200 trials and the experiment duration on average was 36.50 minutes. In every trial, participants will watch one video from the BMP dataset and choose one action from a list of ten options that most accurately describes the action in the video. Videos from all conditions are selected randomly and presented in a random order. All the experiments are conducted with the subjects’ informed consent and according to protocols approved by the Institutional Review Board of our institution. Each subject was compensated.
To control data quality, two checks are implemented in the experiment: (1) Four pre-selected videos in the RGB condition are used as the dummy trials, with two videos representing the "Sit Down" action class and the other two representing the "Stand Up" action class. Since these four videos are meant for quality controls. There is no ambiguity in classifying the action classes in these four videos. These four dummy trials are randomly dispersed with the rest of actual trials in one experiment. Participants who fail to recognize correct actions in any of the four videos are excluded for data analysis. All participants in the experiment passed the dummy trials. (2) Each participant is allowed to participate in the experiment only once. All the trials in one experiment are always unique. This is to prevent the subjects from memorizing the answers from the repeated trials.
Figure 3: Schematic of human psychophysics experiments on Amazon Mechanical Turk (MTurk). In every trial, subjects are presented with a video and a list of ten options. After the video plays only once, the subject has to choose one action among ten options that best describe the action in the video.

### v2

@@ Abstract
Biological motion perception (BMP) refers to humans’ ability to perceive and recognize the actions of living beings solely from their motion patterns, sometimes as minimal as those depicted on point-light displays. While humans excel at these tasks without any prior training, current AI models struggle with poor generalization performance. To close this research gap, we propose the Motion Perceiver (MP). MP solely relies on patch-level optical flows from video clips as inputs. During training, it learns prototypical flow snapshots through a competitive binding mechanism and integrates invariant motion representations to predict action labels for the given video. During inference, we evaluate the generalization ability of all AI models and humans on 62,656 video stimuli spanning 24 BMP conditions using point-light displays in neuroscience. Remarkably, MP outperforms all existing AI models with a maximum improvement of 29% in top-1 action recognition accuracy on these conditions. Moreover, we benchmark all AI models in point-light displays of two standard video datasets in computer vision. MP also demonstrates superior performance in these cases. More interestingly, via psychophysics experiments, we found that MP recognizes biological movements in a way that aligns with human behaviors. Our data and code are available at link.
@@ 1 Introduction
Figure 1: Humans excel at biological motion perception (BMP) tasks with zero training, while current AI models struggle with poor generalization performance. AI models are trained to recognize actions from natural RGB videos and tested using BMP stimuli on point-light displays, which come in two forms: Joint videos, which display only the detected joints of actors in white dots, and Sequential position actor videos (SP), where light points in white are randomly positioned between joints and reallocated to other random positions on the limb in subsequent frames (Sec. 3.1). Note that skeletons, shown in gray in the example video, are not visible to humans or AI models during testing. The generalization performance of both humans and models is assessed after varying five properties in temporal and visual dimensions. See Appendix, Sec. A.1 for example videos.
@@
In parallel to the studies of BMP in psychology and neuroscience, action recognition on images and videos in computer vision has evolved significantly over the past few decades due to its wide range of real-world applications [15, 41, 81, 82, 57, 74]. The field has progressed from relying heavily on hand-crafted features [99, 75, 54] to employing deep-learning-based approaches  [90, 95, 11, 101, 28, 30, 29, 25, 4, 94]. These modern approaches can capture the temporal dynamics and spatial configurations of complex activities within dynamic, unstructured environments  [80, 92]. Despite these advancements, existing AI models still struggle with generalization issues related to occlusion [103, 3, 58], noisy environments  [109], viewpoint variability  [27, 61, 100], subtle human movements  [83, 44] and appearance-free motion information [42]. While various solutions have been proposed to enhance AI generalization in action recognition [12, 72, 14, 70, 104, 48, 59, 113, 79, 50], they do not specifically tackle the generalization challenges in BMP tasks.
@@
Subsequently, to enhance the generalization capability of AI models, we draw inspiration from  [33] in neuroscience and introduce Motion Perceiver (MP). MP only takes dense optical flows between any pairs of video frames as inputs and predicts action labels for the given video. In contrast to many existing pixel-level optical flow models [90, 91, 86, 99], MP calculates dense optical flows at the granularity of patches from the feature maps. In MP, we introduce a set of flow snapshot neurons that learn to recognize and store prototypical motion patterns by competing with one another and binding with dense flows. This process ensures that similar movements activate corresponding snapshot neurons, promoting consistency in motion pattern recognition across patches of video frames. The temporal dynamics within dense flows can vary significantly depending on factors such as the speed, timing, and duration of the actions depicted in video clips. Thus, we also introduce motion-invariant neurons. These neurons decompose motions along four motion directions and integrate their magnitudes over time. This process ensures that features extracted from dense flows remain invariant to small changes and distortions in temporal sequences.
@@ 2.1 Patch-level Optical Flow
Pixel-level optical flow has been a common approach to modelling temporal dynamics in videos  [90, 91, 86, 99]. Different from these works, we use the frozen ViT  [21], pre-trained on ImageNet  [19] with DINO [10], to extract feature map $F_{t}\in\mathbb{R}^{N\times C}$ from $I_{t}$ and compute its optical flow relative to other frames. Consider $F_{t}$ as a 2D grid of patches in $N=H\times W$, where $H$ and $W$ are the height and width of $F_{t}$. $N$ represents the number of patches in the feature map and $C$ is the feature dimension per patch. A unique spatial location for each patch can be defined by its 2D coordinates. We concatenate the X and Y coordinates of all patches in $F_{t}$ to form the patch locations $G\in\mathbb{R}^{N\times 2}$.
@@
Next, we introduce how the patch-level optical flow for video $\mathcal{V}$ is computed. Without loss of generality, given any pair of feature sets $a\in\mathbb{R}^{m\times d}$ and $b\in\mathbb{R}^{n\times d}$, where $m$ and $n$ are the numbers of patches in the feature sets and $d$ is the feature dimension, the adjacency matrix $Q_{a,b}$ between $a$ and $b$ can be calculated as their normalized pairwise feature similarities: $Q_{(a,b)}=\frac{e^{f(a)f(b)^{T}/\tau}}{\sum\limits_{n}e^{f(a)f(b)^{T}/\tau}}\in\mathbb{R}^{m\times n}$, where the superscript $T$ is the transpose function, $f(\cdot)$ is the $l_{2}$-normalization, and $\tau$ is the temperature controlling the sharpness of distribution with its smaller values indicating sharper distribution. We set temperature $\tau=0.001$ and the influence of $\tau$ is analyzed in Appendix, Tab. S4.
@@ 2.2 Flow Snapshot Neurons
Drawing on findings from neuroscience that highlight biological neurons selective for complex optic flow patterns associated with specific movement sequences in motion pathways  [33], we introduce “flow snapshot neurons." These neurons are designed to capture and represent prototypical moments or “slots" in movement sequences. Mathematically, we define $K=6$ flow snapshot neurons or slots $\hat{Z}\in\mathbb{R}^{K\times D}$, where $D=2\times(T-1)$ is the feature dimension per slot. $\hat{Z}$ contains learnable parameters randomly initialized with the Xavier uniform distribution  [35]. The impact of $K$ is discussed in Appendix, Tab. S4.
Slot attention mechanism  [63] separates and organizes different elements of an input into a fixed number of learned prototypical representations or “slots." The learned slots have been useful for semantic segmentation [102, 60, 111, 51, 22]. Here, we follow the training paradigm and implementations of the slot attention mechanism in  [63] and apply it to $\hat{O}$. Specifically, each slot in $\hat{Z}$ attends to unique optical flow sequence patterns in $\hat{O}$ through a competitive attention mechanism based on feature similarities. To ensure the prototypical optical flow patterns in $\hat{Z}$ are diverse, we introduce the loss of contrastive walks $L_{slot}$  [102]. During inference, we keep $\hat{Z}$ frozen and leverage the activations of flow snapshot neurons for action recognition, denoted as $\hat{M}=f(\hat{O})f(\hat{Z})^{T}$, where $\hat{M}\in\mathbb{R}^{T\times N\times K}$. See Appendix, Sec. B.1 and Appendix, Sec. B.2 for mathematical formulations.
Unlike  [108] that use slot-bonded optical flow sequences $\hat{M}\hat{Z}\in\mathbb{R}^{T\times N\times D}$ for downstream tasks, we highlight two advantages of using $\hat{M}$. First, the dimensionality of flow similarities ($K$) is typically much smaller than that of slot-bonded optical flow sequences ($D$), effectively addressing overfitting concerns and reducing computational overhead. Second, by leveraging flow similarities, we benefit from slot activations that encode prototypical flow patterns distilled from raw optical flow sequences. This filtration process helps eliminate noise and irrelevant information from the slot-bonded flow sequences, enhancing the model’s robustness.
@@ 2.3 Motion Invariant Neurons
The temporal dynamics in video clips can vary significantly in the speed and temporal order of the actions. Here, we introduce motion-invariant neurons that capture the accumulative motion magnitudes independent of frame orders. Specifically, we define the optical flow $(O^{x,n,I_{t}}_{m\rightarrow t},O^{y,n,I_{t}}_{m\rightarrow t})$ for patch $n$ in $I_{m}$ moving from frame $I_{m}$ to $I_{t}$ along $x$ and $y$ axes. Every patch-level optical flow in $\hat{O}$ can be projected into four motion components along $+x$, $-x$, $+y$ and $-y$ axes. Without loss of generality, we show that for patch $n$ in $I_{t}$, all its optical flow motions over $T$ frames along the $+x$ axis are aggregated:
@@
By repeating the same operations along $-x,+y,-y$ axes for all the $N$ patches over $T$ frames, we obtain the motion invariant matrix $\tilde{M}\in\mathbb{R}^{T\times N\times 4}$. A stack of self-attention blocks followed by a global average pooling layer fuse information along 4 motion components and then along temporal dimension $T$. To encourage the model to learn motion invariant features for action recognition, we introduce cross-entropy loss $L_{invar}$ to supervise the predicted action against the ground truth $\mathcal{Y}$.
@@ 2.4 Multi-scale Feature Fusion and Training
While $\hat{M}_{1,2,4,8}$ captures the detailed temporal dynamics, $\tilde{M}$ learns robust features against variations in temporal orders. MP combines feature vectors from $\hat{M}_{1,2,4,8}$ and $\tilde{M}$ with a fully connected layer and outputs the final action label. A cross-entropy loss $L_{fuse}$ is used to balance the contributions from $\hat{M}_{1,2,4,8}$ and $\tilde{M}$. The overall loss is: $L=\alpha L_{slot}+L_{flow}+L_{invar}+L_{fuse}$, where the loss weight $\alpha=10$. See Appendix, Tab. S4 for the effect of $\alpha$.
Implementation Details. Our model is trained on Nvidia RTX A5000 and A6000 GPUs, and optimized by AdamW optimizer [65] with cosine annealing scheduler [64] starting from the initial learning rate $10^{-4}$. Data loading is expedited by FFCV [56]. All videos are downsampled to $T=32$ frames. We use a random crop of $224\times 224$ pixels with horizontal flip for training, and a central crop of $224\times 224$ pixels for inference. We use the same set of hyper-parameters for our model in all the datasets. More training details are in Appendix, Sec. C.
@@ 3.1 Our Biological Motion Perception (BMP) Dataset with Human Behavioral Data
Following the works in vision science, psychology, and neuroscience [33, 89, 9, 53, 6, 16, 37, 45, 71, 76, 97], we introduce the BMP dataset, comprising 10 action classes (Appendix, Sec. A.2) specifically chosen for their strong temporal dynamics and minimal reliance on visual cues. Studies [107, 112, 32, 85] indicate that current action recognition models often rely on static features. Point-light displays reduce confounding factors, such as colors, sketches, and body limbs, by minimizing visual information, thereby highlighting the ability to perceive motion. This selection criterion ensures that specific objects or scene contexts, such as a soccer ball on green grass, do not bias the AI models toward recognizing the action as “playing soccer.” This approach aligns with our research focus on generalization in motion perception. In the BMP dataset, there are three types of visual stimuli.
@@
Temporal order (TO): To disrupt the temporal order  [84], we reverse (Reversal) or randomly shuffle (Shuffle) video frames for two types of videos, RGB and Joint above. Temporal resolution (TR): We alter the temporal resolution  [84] by uniformly downsampling 32 frames to 4 or 3 frames, labelled as 4-frames and 3-frames, respectively. For models that need a fixed frame count, each downsampled frame is replicated multiple times before advancing to the next frame in the sequence, until we reach the necessary quantity. Amount of visual information (AVI): We quantify the amount of visual information based on the number of light points (P) in Joint videos  [45]. Specifically, we included conditions: 5, 6, 10, 14, 18, and 26 light points. Lifetime of visual information (LVI): In SP videos, we manipulate the number of consecutive frames during which each point remains at a specific limb position before being randomly reassigned to a different position. Following [6], we refer to this parameter as the “lifetime of visual information" (LT), and include LT values of 1, 2, and 4. A longer lifetime implies that each dot remains at the same limb position for a longer duration, thereby conveying more local image motion information. Invariance to camera views (ICV): Neuroscience research  [43] has revealed brain activity linked to decoding both within-view and cross-view action recognition. To probe view-dependent generalization effects, we sorted video clips from Joints into three categories based on camera views: frontal, $45^{\circ}$, and $90^{\circ}$ views. See Appendix, Sec. A.3 for more implementation details.
Human psychophysics experiments: We conducted human psychophysics experiments schematically illustrated in Appendix, Sec. D, using Amazon Mechanical Turk (MTurk)  [8]. A total of 90 subjects were recruited. For data quality controls, we implemented checks during the experiment. Following a set of filtering criteria, we retained data from 88 subjects. See Appendix, Sec. D for details.